OpenStudy (anonymous):

what is x if x is = to the square root of 2 times 5 square root of 10

OpenStudy (anonymous):

\[\chi =\sqrt{2}\times5\sqrt{10}\]

OpenStudy (anonymous):

So this is like saying sqrt(2)*5*sqrt(10). Multiply the sqrt(2) and sqrt(10) first. You can put them back under the square root together: sqrt(2)*sqrt(10)=sqrt(20)

OpenStudy (anonymous):

so the answer is 5 sqrt 20?

OpenStudy (anonymous):

sqrt(20) can be simplified. 20=4*5